# ESO Expanding Horizon White Paper: Revealing the properties of matter at supranuclear densities with gravitational waves

### v1

@@
Tim Dietrich1,2 , Tanja Hinderer3 , Micaela Oertel4,5, Conrado Albertus Torres6, Nils Andersson7, Dániel Barta8, Andreas Bauswein9, Béatrice Bonga10,11, Marica Branchesi12, G. Fiorella Burgio13, Stefano Burrello14, Prasanta Char6, Sylvain Chaty15, Maria Colonna14, Daniela Doneva16, Anthea F. Fantina17, Tobias Fischer18,19, Juan Garcia-Bellido20, Archisman Ghosh21, Bruno Giacomazzo22,23, Fabian Gittins3, Vanessa Graber24, Francesca Gulminelli25, Jan Harms12,26, Kostas Kokkotas27, Felipe J. Llanes-Estrada28, Michele Maggiore29, Gabriel Martinez-Pinedo9, Andrea Maselli12, Chiranjib Mondal30, Samaya Nissanke1,31, M Ángeles Pérez García6, Cristiano Palomba32, Pantelis Pnigouras33, Anna Puecher1, Michele Punturo34, Adriana R. Raduta35, Violetta Sagun7, Armen Sedrakian18,36, Nikolaos Stergioulas37, Laura Tolos38,39, Kadri Yakut40, Stoytcho Yazadjiev41
@@ 15 Université Paris Cité, CNRS, Astroparticule et Cosmologie, F-75013 Paris, France
18 University of Wroclaw, Wroclaw 50204, Poland
@@ 40 Department of Astronomy and Space Sciences, Faculty of Science, Ege University, 35100, Izmir, Turkey
In general, the increased sensitivity of ET will enable rich multi-messenger observations, with joint GW, electromagnetic (with promising projects such as the eXTP [Li:2025uaw] and SKA [Watts:2014tja]), and neutrino data to deliver a comprehensive view of dense matter across a broad range of physical conditions. Observations by ET and other third-generation GW detectors will unveil details about heavy element synthesis, constrain the EOS across a range of temperatures and compositions, and potentially use NS mergers as probes for exotic physics beyond the Standard Model. With strong support from nuclear theory, ET observations could further be used to test General Relativity and the coupling of gravity to matter in strong-field conditions and investigate potential dark matter effects in compact stars [ET:2025xjr]. Additional unique multi-messenger opportunities with ET for profound insights on dark energy, gravity, and chemical enrichment of the universe are discussed in a companion White Paper about multi-messenger science.

### v2

@@
Tim Dietrich1,2 , Tanja Hinderer3 , Micaela Oertel4,5, Conrado Albertus Torres6, Nils Andersson7, Dániel Barta8, Andreas Bauswein9, Béatrice Bonga10,11, Marica Branchesi12, G. Fiorella Burgio13, Stefano Burrello14, Prasanta Char6, Sylvain Chaty15, Maria Colonna14, Daniela Doneva16, Anthea F. Fantina17, Tobias Fischer18,19, Juan Garcia-Bellido20, Archisman Ghosh21, Bruno Giacomazzo22,23, Fabian Gittins3, Vanessa Graber24, Francesca Gulminelli25, Jan Harms12,26, Kostas Kokkotas27, Felipe J. Llanes-Estrada28, Michele Maggiore29, Gabriel Martinez-Pinedo9, Andrea Maselli12, Chiranjib Mondal30, Samaya Nissanke1,31, M Ángeles Pérez García6, Cristiano Palomba32, Pantelis Pnigouras33, Anna Puecher1, Michele Punturo34, Adriana R. Raduta35, Violetta Sagun7, Armen Sedrakian36,37, Nikolaos Stergioulas38, Laura Tolos39,40, Kadri Yakut41, Stoytcho Yazadjiev42
@@ 15 Université Paris Cité, CNRS, Astroparticule et Cosmologie, F-75013 Paris, France
18 Institute of Theoretical Physics, Wroclaw University of Science and Technology, W. Wyspianskiego 27, 50-370 Wroclaw, Poland
@@ 41 Department of Astronomy and Space Sciences, Faculty of Science, Ege University, 35100, Izmir, Turkey
In general, the increased sensitivity of ET will enable rich multi-messenger observations, with joint GW, electromagnetic (with promising projects such as the eXTP [Li:2025uaw] and SKA [SKAPulsarScienceWorkingGroup:2025ins]), and neutrino data to deliver a comprehensive view of dense matter across a broad range of physical conditions. Observations by ET and other third-generation GW detectors will unveil details about heavy element synthesis, constrain the EOS across a range of temperatures and compositions, and potentially use NS mergers as probes for exotic physics beyond the Standard Model. With strong support from nuclear theory, ET observations could further be used to test General Relativity and the coupling of gravity to matter in strong-field conditions and investigate potential dark matter effects in compact stars [ET:2025xjr]. Additional unique multi-messenger opportunities with ET for profound insights on dark energy, gravity, and chemical enrichment of the universe are discussed in a companion White Paper about multi-messenger science.